OpenStudy (anonymous):

can someone show me how to check my answers for these two problems, first one is 3x^2+5x=0 the answer to that problem is x=-5/3 ....................the second question is 4x^2-13x+3=0 and the answer for that one is x=-1/4 and x=-3

OpenStudy (anonymous):

To check your answer, plug the value you found for x into the equation it's for. If, after simplifying, a true statement emerges (like 0 = 0), then the answer is good. If a false statement emerges (like 1 = 0), then the answer is wrong. 3(-5/3)^2+5(-5/3) = 3*(25/9) - 25/3 = 25/3 - 25/3 = 0 When we plugged the -5/3 in, we got 0 on the left side. This equals 0 on the right side, so -5/3 is a correct answer. 4(-1/4)^2-13(-1/4) + 3 = 4*(1/16)+(13/4) + 3 = 1/4 + 13/4 + 12/4 = 26/4 = 13/2 When we plugged the -1/4 in to the other equation, we got 13/2 on the left side. This does not equal 0 on the right side, so -1/4 is not a correct answer. I think I can leave you to check the x = -3?